a line that passes throught (-6,1) and is parallel to x-15-3y=0
Answer by mananth(16946)

x-3y-15=0
make y the subject
3y=x-15
Find the slope of this line
Divide by -3
y = 1/3 x -5
Compare this equation with y=mx+b
slope m =1/3
The slope of a line parallel to the above line will be the same
The slope of the required line will be 1/3
m= 1/3 ,point ( -6 , 1 )
Find b by plugging the values of m & the point in
y=mx+b
1 = -2.00 + b
b= 3
m= 1/3
Plug value of the slope and b in y = mx +b
The required equation is y = 1/3 x+3
y = 1/ 3 x + 3
m.ananth@hotmail.ca